OpenStudy (anonymous):

10 + x ÷ 2(9 – x^2) for x = -6 solve for x

OpenStudy (anonymous):

pls try to write it through equations

jhonyy9 (jhonyy9):

10+x+18-2x*2=-2x*2+x+28= for x=-6 will be -2(36)-6+28=-72+22=-50